PURCHASE 50 STAMPS. SOME STAMPS .22 EACH. THE REST .14 EACH TOTAL SPENT IS 10.04

Answer by checkley77(12844)

X+Y=50 OR X=50-Y
.22X+.14Y=10.04 REPLACE X BY (50-Y) & SOLVE FOR Y.
.22(50-Y)+.14Y=10.04
11-.22Y+.14Y=10.04
-.08Y=10.04-11
-.08Y=-.96
Y=-.96/-.08
Y=12 ANS.
X+12=50
X=50-12
X=38 ANS.
PROOF:
.22*38+.14*12=10.04
8.36+1.68=10.04
10.04=10.04